Cavy Company estimates that total factory overhead costs will be $777,924 for the year. Direct labor hours are estimated to be 102,900.

a. Compute the predetermined factory overhead rate. Round your answer to the nearest cent.
$ per direct labor hour

b. Determine the amount of factory overhead applied to Job 345 if the amount of direct labor hours is 1,300 and to Job 777 if the amount of direct labor hours is 2,800.

Job 345 $
Job 777 $

c. Journalize the entry to record the factory overhead applied if Jobs 345 and 777 are the only jobs for the period. If an amount box does not require an entry, leave it blank.

Work in Process
Factory Overhead
Answer #1

A) Predetermined overhead rate = Estimated overhead/Estimated direct labor hours

= 777,924/102,900

= 7.56 per dlh

B)

Job 345 = 7.56*1300 = 9828

Job 777 = 7.56*2800 = 21,168

C)

Work in Process 30996
Factory Overhead 30996